### v2

@@ 1. Introduction
In the present article we extend the $J$-invariant to all semisimple algebraic groups. Our idea was that it should be an analogous direct summand $R_{p}(E)$ for groups of outer type whose Poincaré polynomial has cyclotomic type (1.1). Note first of all that it is not true anymore in general that if $G$ is a group of outer type, then the Chow motive of $E/B$ is a direct sum of Tate twists of the same indecomposable motive $R_{p}(G)$. Moreover, our motivic computations showed that the most numerological patterns which occur for groups of inner type, do not apply for Chow motives for groups of outer type.
@@
We show that in both cases the entries of the $J$-invariant of degrees bigger than $2$ do not change when we generically split the respective central simple algebras. Moreover, we explicitly compute the degree one and degree two entries of the $J$-invariant. In the case ${}^{2}\mathrm{A}_{n}$ the degree one entry is related to the discriminant algebra of the respective unitary involution and the degree two entry is related to the underlying central simple algebra. In the case ${}^{2}\mathrm{D}_{n}$ we have an opposite situation, when the degree one entry of the $J$-invariant is related to the underlying central simple algebra and the degree two entry is related to the Clifford algebra of the orthogonal involution (see Proposition 8.2, Corollary 8.10, Proposition 8.15, Corollary 8.19). Analogous statements for groups of inner type were obtained in [Zh22] by the second author.
@@ 3. Preliminaries on the category of Chow motives
For a motive $M$ (resp. for a variety $X$) over a field $F$ and a field extension $E/F$ we denote by $M_{E}$ (resp. by $X_{E}$) the extension of scalars. By $M(n)$ we denote the Tate twist of $M$ by $n$.
Every motive $M$ is determined by a smooth projective variety $X$ and a projector ${\pi\in\mathop{\mathrm{Ch}}\nolimits(X\times X)}$. A motivic decomposition of a motive $M$ is a decomposition into a direct sum. Motivic decompositions of a motive $M=(X,\pi)$ correspond to decompositions of the projector $\pi$ into a sum of (pairwise) orthogonal projectors. The realization (resp. the co-realization) of $M=(X,\pi)$ is the group $(p_{2})_{*}(\pi\cdot p_{1}^{*}(\mathop{\mathrm{Ch}}\nolimits(X)))$ (resp. $(p_{1})_{*}(\pi\cdot p_{2}^{*}(\mathop{\mathrm{Ch}}\nolimits(X)))$), where $p_{i}\colon X\times X\to X$ is the projection to the $i$-th factor ($i=1,2$), $p_{i}^{*}$ is the pullback, $(p_{i})_{*}$ is the push-forward and $\cdot$ is the intersection product. By $\circ$ we denote the composition product.
@@
Let $X$ and $Y$ be smooth projective equidimensional varieties over $F$. A correspondence $\alpha\colon X\rightsquigarrow Y$ of degree $l$ is an element in $\mathop{\mathrm{Ch}}\nolimits_{\dim X+l}(X\times Y)$. The multiplicity of $\alpha$ is an integer $\operatorname{mult}(\alpha)$ such that $p_{*}(\alpha)=\operatorname{mult}(\alpha)\cdot[X]$, where $p_{*}$ is the push-forward of the projection $p\colon X\times Y\to X$. By $\operatorname{Corr}(X,Y)$ we denote the group of all correspondences from $X$ to $Y$ over $F$ of degree 00. We also write $\operatorname{End}_{\operatorname{CM}}(X)=\operatorname{Corr}(X,X)$. This is the group of endomorphisms of $M(X)$ in the category of motives. If $A$ is a matrix consisting of correspondences, then we apply $\operatorname{mult}$ to $A$ elementwise (in particular, $\operatorname{mult}A$ is a matrix of the same size as $A$).
@@
and for a Tate motive $\mathbb{T}(i)$ we have
(3.4) $R_{L/F}(\mathbb{T}(i))=\mathbb{T}(mi)$
@@
Besides, for a smooth variety $X$ over $L$ we will use a natural map
@@ 4. Background on the $J$-invariant for groups of inner type
Let $\mathrm{B}B$ denote the classifying space of $B$. By [Gr58, (4.1)] there is a characteristic map
@@
In fact, one can replace $B$ in sequence (4.3) by any special parabolic subgroup $P$ of $G$ (see [PS17, Lemma 7.1]). Note also that by [KM06, Theorem 6.4] the image of $c$ coincides with the subring of rational cycles for the twisted form (${}_{E}(G/B)$ (or, more generally, (${}_{E}(G/P)$ for a special parabolic subgroup $P$) under a generic $G$-torsor $E$.
[…]
@@
[…]
@@
$\operatorname{\mathrm{SL}}_{n}/\mu_{m},\,m|n$ $p|m$ $1$ $1$ $p^{k_{1}}\parallel n$
[…]
$\mathrm{O}^{+}_{n}$, $n>2$ $2$ $[\frac{n+1}{4}]$ $12i-$ 1 $[\log_{2}\frac{n-1}{2i-1}]$ if $d_{i}+l=2^{s}d_{m}$ and
@@
[…]
$2i-1,\,i\geq 2$ $[\log_{2}\frac{2n-1}{2i-1}]$
$\operatorname{\mathrm{Spin}}_{n}$, $n>2$ $2$ $[\frac{n-3}{4}]$ $12i+$ 1 $[\log_{2}\frac{n-1}{2i+1}]$ the same restrictions
$\operatorname{\mathrm{PGO}}_{2n}^{+}$, $n>1$ $2$ $[\frac{n+2}{2}]$ $11,\,i=$ 1 $2^{k_{1}}\parallel n$ the same restrictions
$2i-3,\,i\geq 2$ $[\log_{2}\frac{2n-1}{2i-3}]$ assuming $i,m\geq 2$
$\mathrm{G}_{2}$ $2$ $1$ $3$ $1$
$\mathrm{F}_{4}$ $2$ $1$ $3$ $1$
$\mathrm{E}_{6}$ $2$ $1$ $3$ $1$
$\mathrm{F}_{4}$ $3$ $1$ $4$ $1$
$\mathrm{E}_{6}^{sc}$ $3$ $1$ $4$ $1$
$\mathrm{E}_{7}$ $3$ $1$ $4$ $1$
$\mathrm{E}_{6}^{ad}$ $3$ $2$ 1, 4141,\,41 , 4 2, 1212,\,12 , 1 $|j_{1}-j_{2}|\leq 1$
$\mathrm{E}_{7}^{sc}$ $2$ $3$ 3, 5, 93593,\,5,\,93 , 5 , 9 1, 1, $,\,1,\,1$ 1 $j_{1}\geq j_{2}\geq j_{3}$
$\mathrm{E}_{7}^{ad}$ $2$ $4$ 1, 3, 5, 913591,\,3,\,5,\,91 , 3 , 5 , 9 1, 1, 1, $,\,1,\,1,\,1$ 1 $j_{2}\geq j_{3}\geq j_{4}$
$\mathrm{E}_{8}$ $2$ $4$ 3, 5, 9, 15359153,\,5,\,9,\,153 , 5 , 9 , 15 3, 2, 1, 132113,\,2,\,1,\,13 , 2 , 1 , 1 $j_{1}\geq j_{2}\geq j_{3}$,
@@ 5. Normed Chow rings and motives
From now on we assume that $[K:F]=p^{n}$ for some prime number $p$ and for some $n\geq 1$, since otherwise the respective normed Chow rings are zero. Note also that then $\mathop{\mathrm{CH}}\nolimits_{K}(X)$ is an algebra over $\mathop{\mathrm{CH}}\nolimits_{K}(\mathrm{pt})=\mathbb{F}_{p}$ where $\mathrm{pt}$ denotes $\operatorname{Spec}F$. Furthermore, for a smooth variety $Y$ over $E$ we have ${\mathop{\mathrm{CH}}\nolimits_{K}(\mathop{\mathrm{cor}}\nolimits_{E/F}(Y))=0}$ where ${F\subsetneq E\subset K}$.
@@
Moreover, by [KM06, Theorem 6.4] the image of $c$ coincides with the subring of rational cycles for the twisted form (${}_{E}(G^{qs}/B^{qs})$ under a generic $G^{qs}$-torsor $E$ over $F$.
Furthermore, we have for an adjoint or simply connected group $G^{qs}$ that ${\mathop{\mathrm{CH}}\nolimits^{*}(\mathrm{B}B^{qs})=\mathop{\mathrm{CH}}\nolimits^{*}(\mathrm{B}B)^{\Gamma}}$ and $\mathop{\mathrm{CH}}\nolimits^{*}(G^{qs}/B^{qs})=\mathop{\mathrm{CH}}\nolimits^{*}(G/B)^{\Gamma}$, where $\Gamma$ denotes the absolute Galois group, and $B$ and $G$ denote $B^{qs}$ and $G^{qs}$ extended to a separable closure of $F$.
Passing to the normed Chow rings we obtain a characteristic map
@@
[…]
@@
If $X$ is a twisted flag variety, then we define its normed upper motive $\mathcal{U}_{K}(X)$ exactly in the same way as its ordinary upper motive (see Section 3), but replacing the ordinary Chow groups by the normed Chow groups.
Note that in the context of normed Chow rings we will denote by $\overline{X}$ (and, more generally, by $\overline{N}$ for a direct summand $N$ of the normed motive of $X$) the variety $X$ not over a splitting field of the respective group, but over the function field of a generically quasi-split variety (e.g., the variety of Borel subgroups). Over such fields the respective group is quasi-split, but the field $K$ remains non-split. In particular, for a normed motive $N$ we write $P(\mathop{\mathrm{CH}}\nolimits_{K}^{*}(N),t)$ for the Poincaré polynomial of $N$ over such quasi-splitting field.
@@ 6. Some numerology of algebraic groups
$\mathrm{B}_{m},\mathrm{C}_{m},\mathrm{D}_{m},\mathrm{G}_{2}$ $2$
@@
${}^{2}\mathrm{A}_{m},{}^{2}\mathrm{D}_{m},{}^{2}\mathrm{E}_{6}$ $2$
${}^{3}\mathrm{D}_{4}$ $3$
@@
${}^{2}\mathrm{D}_{m}$ $\mathrm{B}_{m-1}$ $22,4,\ldots,2m-$ 2 $m$
@@
d) Let $G=R_{E/F}(G^{\prime})$ for a non-trivial finite separable field extension $E/F$ and an absolutely simple algebraic quasi-split group $G^{\prime}$ over $E$. Assume that $[K:F]$ is a power a prime number $p$. Let $B^{\prime}$ be a Borel subgroup of $G^{\prime}$ defined over $E$ and $m=[E:F]$.
@@
Part d) follows from the properties of the Weil restrictions summarized in Section 3 and the fact that the normed Chow ring eliminates all cycles from the ordinary Chow ring which are not Galois invariant (see Section 5). ∎
@@
Then the group $G$ is of even orthogonal type and the variety $Q=G/P_{1}$ is a (quasi-split) projective quadric of dimension $22n-$ 2.
@@ 7. Motives of generically quasi-split varieties
First of all, we need to replace $\mathop{\mathrm{Ch}}\nolimits^{*}(G)$ in Definition 4.6 by the ring $C^{*}$ from sequence (5.3). Moreover, we need to know that as in the inner case this ring has presentation (4.5) and we need to compute the parameters $r$, $d_{i}$, and $k_{i}$ for this presentation. Once this is done the same proof as in the inner case works and we obtain an analogous motivic decomposition for normed motives as in the inner case (see Theorem 8.21 below).
Let $Y$ be a generically quasi-split twisted flag variety. We denote by $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(Y)$ the subring of rational cycles on $Y$ over $F(Y)$. In order to describe the ring $C^{*}$ we need to investigate the structure of the ring $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ as a $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(Y)$-module. To begin with we have the following well-known fact about the Chow group of the upper motive of $Y$.
@@
Indeed, since $Y$ is generically quasi-split, every rational cycle in $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\mathcal{U}_{K}(Y))$ is a generic point of some direct summand of $\mathcal{U}_{K}(Y)$. The claim follows now, since the motive $\mathcal{U}_{K}(Y)$ is indecomposable. ∎
We fix now a projector $\pi\in\operatorname{End}(M_{K}(Y))$ of the normed motive of $Y$ such that $(Y,\pi)\simeq\mathcal{U}_{K}(Y)$. Recall that $\operatorname{End}(Y,\pi)=\pi\circ\operatorname{End}(M_{K}(Y))\circ\pi\subset\mathop{\mathrm{CH}}\nolimits_{K}^{\dim Y}(Y\times Y)$. Denote by $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))$ the image of the restriction map ${\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))\rightarrow{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(\overline{Y,\pi}))$ and write $\overline{\pi}$ for the image of $\pi$ in $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))$.
@@
$\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))=\mathbb{F}_{p}\cdot\overline{\pi}$.
@@
Note that, since $(Y,\pi)\simeq\mathcal{U}_{K}(Y)$, we have $\operatorname{mult}\overline{\pi}=1$. Let $\overline{\alpha}\in\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))$ and let $\alpha$ be its preimage in ${\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))$. Replacing if necessary $\alpha$ by $\alpha-\lambda\cdot\pi$ with $\lambda\in\mathbb{F}_{p}$, we can assume that $\operatorname{mult}\alpha=0$ and our goal is to show that $\alpha=0$.
We fix a homogeneous $\mathbb{F}_{p}$-basis $\mathcal{B}$ of $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ with its dual basis $\mathcal{B}^{*}$ and write $\overline{\alpha}=\operatornamewithlimits{\textstyle\sum}_{i\in I}a_{i}\times b_{i}$ for some index set $I$, where $b_{i}\in\mathcal{B}$. We choose $a_{l}$, $l\in I$, with the minimal codimension among all elements $a_{i},i\in I$. Note that, since $\operatorname{mult}\alpha=0$, we have $\operatorname{codim}a_{l}>0$. Consider the cycle $b_{l}^{*}\in\mathcal{B}^{*}$ dual to $b_{l}$. That is, for every $i\in I$ the degree $\deg(b_{i}b_{l}^{*})$ is equal to $1$ if $i=l$ and is equal to 00 otherwise. Since the normed motive of $Y$ splits over the function field $F(Y)$, we can identify $\mathop{\mathrm{CH}}\nolimits_{K}(Y_{F(Y)})$ with $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$.
@@
Since $\overline{\beta}=1\times b_{l}^{*}+\dots$, where ``"``"`$\dots$"` ` … " stands for a linear combination of only those terms whose first factor has codimension $>0$, we have
@@
where $[\mathrm{pt}]$ denotes the class of a rational point and ``"``"`$\dots$"` ` … " stands for a linear combination of only those terms whose second factor has dimension $>0$. It follows that $a_{l}$ is $F$-rational. Since $a_{l}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$ and $\operatorname{codim}a_{l}>0$, by Lemma 7.1 we have $a_{l}=0$ and, therefore, $\alpha=0$. ∎
@@
Let $\overline{\alpha}\in\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(\operatorname{End}(Y,\pi))$. Then

### v3

@@ 1. Introduction
In the present article we extend the $J$-invariant to all semisimple algebraic groups. Our idea was that it should be an analogous direct summand $R_{p}(E)$ for groups of outer type whose Poincaré polynomial has cyclotomic type (1.1). Note first of all that it is not true anymore in general that if $G$ is a group of outer type, then the Chow motive of $E/B$ is a direct sum of Tate twists of the same indecomposable motive $R_{p}(G)$. Moreover, our motivic computations showed that the most numerological patterns, which occur for groups of inner type, do not apply for Chow motives for groups of outer type.
@@
We show that in both cases the entries of the $J$-invariant of degrees bigger than $2$ do not change when we generically split the respective central simple algebras. Moreover, we explicitly compute the degree one and degree two entries of the $J$-invariant. In the case $^{2}\mathrm{A}_{n}$ the degree one entry is related to the discriminant algebra of the respective unitary involution and the degree two entry is related to the underlying central simple algebra. In the case $^{2}\mathrm{D}_{n}$ we have an opposite situation, when the degree one entry of the $J$-invariant is related to the underlying central simple algebra and the degree two entry is related to the Clifford algebra of the orthogonal involution (see Proposition 8.2, Corollary 8.10, Proposition 8.15, Corollary 8.19). Analogous statements for groups of inner type were obtained in [Zh24] by the second author.
@@ 3. Preliminaries on the category of Chow motives
For a motive $M$ (resp. for a variety $X$) over a field $F$ and a field extension $E/F$ we denote by $M_{E}$ (resp. by $X_{E}$) the extension of scalars. The motive of the projective line $\mathbb{P}^{1}$ over $F$ decomposes into a direct sum $\mathbb{F}_{p}\oplus\mathbb{F}_{p}(1)$, where $\mathbb{F}_{p}$ denotes the motive of $\operatorname{Spec}F$ and $\mathbb{F}_{p}(1)$ is the complementary direct summand called the Tate motive. By $M(n):=M\otimes\mathbb{F}_{p}(1)^{\otimes n}$ we denote the Tate twist of $M$ by $n$.
Every motive $M$ is determined by a smooth projective variety $X$ and a projector $\pi\in\mathop{\mathrm{Ch}}\nolimits(X\times X)$}. A motivic decomposition of a motive $M$ is a decomposition into a direct sum. Motivic decompositions of a motive $M=(X,\pi)$ correspond to decompositions of the projector $\pi$ into a sum of (pairwise) orthogonal projectors. The realization (resp. the co-realization) of a cycle $\beta\in\mathop{\mathrm{Ch}}\nolimits(X\times X)$ is the map $\beta_{*}\colon\mathop{\mathrm{Ch}}\nolimits(X)\to\mathop{\mathrm{Ch}}\nolimits(X)$ (resp. $\beta^{*}\colon\mathop{\mathrm{Ch}}\nolimits(X)\to\mathop{\mathrm{Ch}}\nolimits(X)$) given by $x\mapsto(p_{2})_{*}(\pi\cdot p_{1}^{*}(x))$ (resp. $x\mapsto(p_{1})_{*}(\pi\cdot p_{2}^{*}(x))$), where $p_{i}\colon X\times X\to X$ is the projection to the $i$-th factor ($i=1,2$), $p_{i}^{*}$ is the pullback, $(p_{i})_{*}$ is the push-forward and $\cdot$ is the intersection product. By $\circ$ we denote the composition product.
@@
Let $X$ and $Y$ be smooth projective equidimensional varieties over $F$. A correspondence $\alpha\colon X\rightsquigarrow Y$ of degree $l$ is an element in $\mathop{\mathrm{Ch}}\nolimits_{\dim X+l}(X\times Y)$. The multiplicity of $\alpha$ is an element $\operatorname{mult}(\alpha)\in\mathbb{F}_{p}$ such that $\operatorname{\mathrm{pr}}_{*}(\alpha)=\operatorname{mult}(\alpha)\cdot[X]$, where $\operatorname{\mathrm{pr}}_{*}$ is the push-forward of the projection $\operatorname{\mathrm{pr}}\colon X\times Y\to X$. By $\operatorname{Corr}(X,Y)$ we denote the group of all correspondences from $X$ to $Y$ over $F$ of degree 0. We also write $\operatorname{End}_{\operatorname{CM}}(X)=\operatorname{Corr}(X,X)$. This is the group of endomorphisms of $M(X)$ in the category of motives.
@@
and for a Tate motive $\mathbb{F}_{p}(i)$ we have
(3.4) $R_{L/F}(\mathbb{F}_{p}(i))=\mathbb{F}_{p}(mi)$
@@
Besides, for a smooth variety $X$ over $L$ we will use in Section 8 a natural map
@@ 4. Background on the $J$-invariant for groups of inner type
Let $\mathrm{B}B$ denote the classifying space of $B$ (see [EG98], [To99]). By [Gr58, (4.1)] there is a characteristic map
@@
(4.3) $\mathop{\mathrm{CH}}\nolimits^{*}(\mathrm{B}B)\xrightarrow{c}\mathop{\mathrm{CH}}\nolimits^{*}(G/B)\xrightarrow{\pi}\mathop{\mathrm{CH}}\nolimits^{*}(G)$
@@
4.4JJ-invariant for Chow motives.
@@
$\mathop{\mathrm{Ch}}\nolimits^{*}(E/B)\xrightarrow{\mathop{\mathrm{res}}\nolimits}\mathop{\mathrm{Ch}}\nolimits^{*}(G/B)\xrightarrow{\pi}\mathop{\mathrm{Ch}}\nolimits^{*}(G),$
@@
$\operatorname{\mathrm{SL}}_{n}/\mu_{m},\,m|n$ $p|m$ 11 11 $p^{k_{1}}\parallel n$
$\operatorname{\mathrm{PGSp}}_{n},\,2|n$ 22 11 11 $2^{k_{1}}\parallel n$
$\mathrm{O}^{+}_{n}$, $n>2$ 22 $\frac{n+1}{4}]$ 2i12i-1 $\log_{2}\frac{n-1}{2i-1}]$ if $d_{i}+l=2^{s}d_{m}$ and
@@
$\operatorname{\mathrm{Spin}}_{2n}^{\pm},\,2|n$ 22 $\frac{n}{2}$ 1,i=11,\,i=1 $2^{k_{1}}\parallel n$ the same restrictions
$2i-1,\,i\geq 2$ $\log_{2}\frac{2n-1}{2i-1}]$
$\operatorname{\mathrm{Spin}}_{n}$, $n>2$ 22 $\frac{n-3}{4}]$ $2i+1$ $\log_{2}\frac{n-1}{2i+1}]$ the same restrictions
$\operatorname{\mathrm{PGO}}_{2n}^{+}$, $n>1$ 22 $\frac{n+2}{2}]$ 1,i=11,\,i=1 $2^{k_{1}}\parallel n$ the same restrictions
$2i-3,\,i\geq 2$ $\log_{2}\frac{2n-1}{2i-3}]$ assuming $i,m\geq 2$
$\mathrm{G}_{2}$ 22 11 33 11
$\mathrm{F}_{4}$ 22 11 33 11
$\mathrm{E}_{6}$ 22 11 33 11
$\mathrm{F}_{4}$ 33 11 44 11
$\mathrm{E}_{6}^{sc}$ 33 11 44 11
$\mathrm{E}_{7}$ 33 11 44 11
$\mathrm{E}_{6}^{ad}$ 33 22 1, 41,\,4 2, 12,\,1 $|j_{1}-j_{2}|\leq 1$
$\mathrm{E}_{7}^{sc}$ 22 33 3, 5, 93,\,5,\,9 1, 1, 11,\,1,\,1 $j_{1}\geq j_{2}\geq j_{3}$
$\mathrm{E}_{7}^{ad}$ 22 44 1, 3, 5, 91,\,3,\,5,\,9 1, 1, 1, 11,\,1,\,1,\,1 $j_{2}\geq j_{3}\geq j_{4}$
$\mathrm{E}_{8}$ 22 44 3, 5, 9, 153,\,5,\,9,\,15 3, 2, 1, 13,\,2,\,1,\,1 $j_{1}\geq j_{2}\geq j_{3}$,
@@ 5. Normed Chow rings and motives
From now on we assume that $K:F]=p^{n}$ for some prime number $p$ and for some $n\geq 1$, since otherwise the respective normed Chow rings are zero. Note also that then $\mathop{\mathrm{CH}}\nolimits_{K}(X)$ is an algebra over $\mathop{\mathrm{CH}}\nolimits_{K}(\mathrm{pt})=\mathbb{F}_{p}$ where $\mathrm{pt}$ denotes $\operatorname{Spec}F$. Furthermore, for a smooth variety $Y$ over $E$ we have $\mathop{\mathrm{CH}}\nolimits_{K}(\mathop{\mathrm{cor}}\nolimits_{E/F}(Y))=0$} where $F\subsetneq E\subset K$}.
@@
[…]
Furthermore, we have for an adjoint or simply connected group $G^{qs}$ that
$\mathop{\mathrm{CH}}\nolimits^{*}(\mathrm{B}B^{qs})=\mathop{\mathrm{CH}}\nolimits^{*}(\mathrm{B}B)^{\Gamma}$}
and $\mathop{\mathrm{CH}}\nolimits^{*}(G^{qs}/B^{qs})=\mathop{\mathrm{CH}}\nolimits^{*}(G/B)^{\Gamma}$, where $\Gamma$ denotes the absolute Galois group, and $B$ and $G$ denote $B^{qs}$ and $G^{qs}$ extended to a separable closure of $F$.
Let $K$ be a Galois field extension of $F$ of $p$-primary degree such that $G^{qs}_{K}$ is a group of inner type. Passing to the normed Chow rings we obtain a characteristic map
@@
(5.3) $\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\mathrm{B}B^{qs})\xrightarrow{c}\mathop{\mathrm{CH}}\nolimits_{K}^{*}(G^{qs}/B^{qs})\to C^{*}$
@@
Let $X$ be a twisted flag variety over $F$ and let $K$ be a Galois field extension of $F$ of $p$-primary degree that that the respective group is a group of inner type over $K$. We define the normed upper motive $\mathcal{U}_{K}(X)$ of $X$ exactly in the same way as its ordinary upper motive (see Section 3), but replacing the ordinary Chow groups by the normed Chow groups, i.e., $\mathcal{U}_{K}(X)$ is an indecomposable direct summand of the normed motive $M_{K}(X)$ of $X$ such that $\mathop{\mathrm{CH}}\nolimits_{K}^{0}(\mathcal{U}_{K}(X))\neq 0$.
Note that in the context of normed Chow rings we will denote by $\overline{X}$ (and, more generally, by $\overline{N}$ for a direct summand $N$ of the normed motive of $X$) the variety $X$ not over a splitting field of the respective group, but over the function field $L$ of a generically quasi-split variety (e.g., the variety of Borel subgroups). Over such fields the respective group is quasi-split, but the field $K$ remains a field, i.e., $K\otimes_{F}L$ is still a field.
Note that in the category of normed Chow motives every direct summand of a quasi-split twisted flag variety is a direct sum of (normed) Tate motives. Therefore, for a normed motive $N$ as in the previous paragraph we can define its normed Poincaré polynomial $P(\mathop{\mathrm{CH}}\nolimits_{K}^{*}(N),t)$ as $\sum_{i\geq 0}\dim\mathop{\mathrm{CH}}\nolimits_{K}^{i}(\overline{N})\,t^{i}$, where the motive $\overline{N}$ is considered over a quasi-splitting field as above.
@@ 6. Some numerology of algebraic groups
$\mathrm{B}_{m},\mathrm{C}_{m},\mathrm{D}_{m},\mathrm{G}_{2}$ 22
@@
$^{2}\mathrm{A}_{m},{}^{2}\mathrm{D}_{m},{}^{2}\mathrm{E}_{6}$ 22
$^{3}\mathrm{D}_{4}$ 33
@@
$^{2}\mathrm{D}_{m}$ $\mathrm{B}_{m-1}$ $2,4,\ldots,2m-2$ $m$
@@
d) Let $G=R_{E/F}(G^{\prime})$ for a non-trivial finite separable field extension $E/F$ and an absolutely simple algebraic quasi-split group $G^{\prime}$ over $E$. Assume that $[K:F]$ is a power of a prime number $p$. Let $B^{\prime}$ be a Borel subgroup of $G^{\prime}$ defined over $E$ and $m=[E:F]$.
@@
Part d) follows from the properties of the Weil restrictions summarized in Section 3 and the fact that the normed Chow ring eliminates all cycles from the ordinary Chow ring which are not Galois invariant (see Section 5). Note also that $K$ contains $E$, since by the assumptions the group $G$ is of inner type over $K$. ∎
@@
Then the group $G$ is of even orthogonal type and the variety $Q=G/P_{1}$ is a (quasi-split) projective quadric of dimension 2n22n-2.
@@ 7. Motives of generically quasi-split varieties
First of all, we need to replace $\mathop{\mathrm{Ch}}\nolimits^{*}(G)$ in Definition 4.6 by the ring $C^{*}$ from sequence (5.3). Moreover, we need to know that, as in the inner case, this ring has presentation (4.5) and we need to compute the parameters $r$, $d_{i}$, and $k_{i}$ for this presentation. Once this is done, the same proof as in the inner case works, and we obtain an analogous motivic decomposition for normed motives as in the inner case (see Theorem 8.21 below).
Let $Y$ be a generically quasi-split twisted flag variety over $F$. Let $K$ be a Galois field extension of $F$ of $p$-primary degree such that the respective group is a group of inner type over $K$. We denote by $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(Y)$ the subring of rational cycles on $Y$ over $F(Y)$. In order to describe the ring $C^{*}$ we need to investigate the structure of the ring $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ as a $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(Y)$-module, where $\overline{Y}$ stands for $Y_{F(Y)}$. Note that since $Y$ is generically quasi-split, the normed motive of $Y$ over $F(Y)$ is a direct sum of Tate motives.
To begin with, we have the following well-known fact about the Chow group of the upper motive of $Y$.
@@
Since $Y$ is generically quasi-split, every rational cycle in ${\mathop{\mathrm{CH}}\nolimits}_{K}(\mathcal{U}_{K}(Y))$ is a generic point of some direct summand of $\mathcal{U}_{K}(Y)$ (the proof is the same as for the inner case; see [GPS16, Corollary 4.11]).
The lemma now follows, since the motive $\mathcal{U}_{K}(Y)$ is indecomposable. ∎
We fix now a projector $\pi\in\operatorname{End}(M_{K}(Y))$ of the normed motive of $Y$ such that $(Y,\pi)\simeq\mathcal{U}_{K}(Y)$. Recall that $\operatorname{End}(Y,\pi)=\pi\circ\operatorname{End}(M_{K}(Y))\circ\pi\subset\mathop{\mathrm{CH}}\nolimits_{K}^{\dim Y}(Y\times Y)$. Denote by $\overline{\operatorname{End}}(Y,\pi)$ the image of the restriction map $\operatorname{End}(Y,\pi)\rightarrow\operatorname{End}(\overline{Y,\pi})$ and write $\overline{\pi}$ for the image of $\pi$ in $\overline{\operatorname{End}}(Y,\pi)$.
For cycles in the normed Chow groups we consider their multiplicities $\operatorname{mult}$ defined exactly in the same way as in the case of ordinary Chow groups (see Section 3).
@@
$\overline{\operatorname{End}}(Y,\pi)=\mathbb{F}_{p}\cdot\overline{\pi}$.
@@
Note that, since $(Y,\pi)\simeq\mathcal{U}_{K}(Y)$, we have $\operatorname{mult}\overline{\pi}=1$. Let $\overline{\alpha}\in\overline{\operatorname{End}}(Y,\pi)$ and let $\alpha$ be its preimage in $\operatorname{End}(Y,\pi)$. Replacing if necessary $\alpha$ by $\alpha-\lambda\cdot\pi$ with $\lambda\in\mathbb{F}_{p}$, we can assume that $\operatorname{mult}\alpha=0$ and our goal is to show that $\alpha=0$.
We fix a homogeneous $\mathbb{F}_{p}$-basis $\mathcal{B}$ of $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ with its dual basis $\mathcal{B}^{*}$ and write
$\overline{\alpha}=\operatornamewithlimits{\textstyle\sum}_{i\in I}a_{i}\times b_{i}$}
for some index set $I$, where $b_{i}\in\mathcal{B}$. We choose $a_{l}$, $l\in I$, with the minimal codimension among all elements $a_{i},i\in I$. Note that, since $\operatorname{mult}\alpha=0$, we have $\operatorname{codim}a_{l}>0$}. Consider the cycle $b_{l}^{*}\in\mathcal{B}^{*}$ dual to $b_{l}$. That is, for every $i\in I$ the degree $\deg(b_{i}b_{l}^{*})$ is equal to $1$ if $i=l$ and is equal to 0 otherwise.
@@
Since $\overline{\beta}=1\times b_{l}^{*}+\dots$, where $\dots$" stands for a linear combination of only those terms whose first factor has codimension >0>0, we have
@@
where $\mathrm{pt}]$ denotes the class of a rational point and $\dots$" stands for a linear combination of only those terms whose second factor has dimension >0>0. It follows that $a_{l}$ is $F$-rational. Since $a_{l}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$ and $\operatorname{codim}a_{l}>0$, by Lemma 7.1 we have $a_{l}=0$ and, therefore, $\alpha=0$. ∎
@@
Let $\overline{\alpha}\in\overline{\operatorname{End}}(Y,\pi)$. Then